# Cyber-Twin: Digital Twin-boosted Autonomous Attack Detection for Vehicular Ad-Hoc Networks

### 09 Feb 2024 v3

@@ Abstract
The rapid evolution of Vehicular Ad-hoc NETworks (VANETs) has ushered in a transformative era for intelligent transportation systems (ITS), significantly enhancing road safety and vehicular communication. However, the intricate and dynamic nature of VANETs presents formidable challenges, particularly in vehicle-to-infrastructure (V2I) communications. Roadside Units (RSUs), integral components of VANETs, are increasingly susceptible to cyberattacks, such as jamming and distributed denial-of-service (DDoS) attacks. These vulnerabilities pose grave risks to road safety, potentially leading to traffic congestion and vehicle malfunctions. Current approaches often struggle to effectively merge digital twin technology with Artificial Intelligence (AI) models to boost security and sustainability. In response to this gap, our study introduces an innovative cyber-twin framework tailored to enhance the security of RSUs in VANETs. This framework uniquely combines digital twin technology with cutting-edge AI to offer a real-time, dynamic representation of RSUs. This allows for detailed monitoring and efficient detection of threats, significantly strengthening RSU security in VANETs. Moreover, our framework makes a notable contribution to eco-friendly communication by improving the computational efficiency of RSUs, leading to increased energy efficiency and extended hardware durability. Our results show a considerable enhancement in resource management and attack detection, surpassing the performance of existing solutions. In particular, the cyber-twin framework showed a substantial reduction in RSU load and an optimal balance between resource consumption and high attack detection efficiency, with a defined twinning rate range of seventy-six to ninety per cent. These advancements underscore our commitment to developing sustainable, secure, and resilient vehicular communication systems for the future of smart cities.
@@ I Introduction
In recent years, the advent of Vehicular Ad-hoc NETworks (VANETs) has marked a significant milestone in the evolution of intelligent transportation systems (ITS). These networks, which seamlessly integrate mobile ad hoc networks (MANET), the Internet of Things (IoT), and intelligent transport systems, are at the forefront of revolutionizing road safety and vehicular communication. The critical role of VANETs in reducing traffic accidents and enhancing road safety cannot be overstated, given the alarming statistics from the World Health Organization highlighting the high incidence of road traffic deaths globally [1].
@@
However, the dynamic and complex nature of VANETs introduces significant challenges, particularly in vehicle-to-infrastructure communications [2]. Road Side Units (RSUs), as critical components of VANETs, are often the targets of cyberattacks, including but not limited to jamming and distributed denial-of-service (DDoS) attacks, which can severely disrupt vehicular communication and pose significant risks to road safety, leading to potential traffic congestion, vehicle malfunctions, and even accidents [3]. This vulnerability underscores the urgency for innovative solutions to safeguard these systems against such threats. The urgent need to address these vulnerabilities is underscored by the rapid evolution of vehicular technology and the increasing prevalence of connected and autonomous vehicles in modern urban environments.
Figure 1: The cyber-twin architecture.
@@
In addressing these challenges, our work focuses primarily on vehicle-to-infrastructure communication, explicitly targeting the security of RSUs. We introduce a novel cyber-twin framework that employs digital twin technology and an advanced Artificial Intelligence (AI) model. The digital twin concept, a digital replica of a physical system, is central to our approach [4]. By creating a digital twin of RSUs, we achieve a real-time, dynamic representation of the physical units, which allows for the meticulous monitoring and analysis of network interactions. This capability is crucial in detecting and mitigating the aforementioned cyber threats, ensuring the integrity and reliability of vehicle-to-infrastructure communications [5] Furthermore, the significance of our solution lies not only in its security aspects but also in its contribution to green communication. By optimizing the efficiency of RSUs and reducing their computational demands, our framework aligns with the growing emphasis on sustainable and eco-friendly technologies in smart cities. This dual focus on security and sustainability positions our work at the forefront of advancing intelligent transportation systems. Moreover, this reduction in energy consumption and prolongation of hardware lifespans underscores our commitment to developing environmentally conscious technologies for smart cities. The key contributions of this paper include:
@@
  • We propose a cyber-twin framework tailored explicitly for securing vehicle-to-infrastructure communication in VANETs, utilizing an AI model for efficient threat detection.
@@
  • By integrating digital twin technology, our approach provides a novel and effective method for real-time monitoring, management of vehicular networks, and analysis of RSU performance and security.
@@
  • The cyber-twin framework significantly reduces the computational load on RSUs, contributing to the longevity of hardware and overall energy efficiency.
  • Our solution has significant advancements in green communication practices, contributing to the sustainability of intelligent transportation systems.
@@
This paper presents a pioneering approach to enhancing the security and efficiency of vehicle-to-infrastructure communication in VANETs while simultaneously addressing the critical aspects of road safety and environmental sustainability. We aim to set a new standard in developing intelligent and secure transportation systems through the novel application of digital twin technology and AI models. In essence, our work addresses a critical gap in vehicular network security. It paves the way for more sustainable and intelligent transportation systems, marking a significant stride in the journey towards safer and greener smart cities. This work proceeds with an overview of relevant literature in Section II, then delves into the proposed methodology in Section III, and evaluates its performance in Section IV. The paper is drawn to a close with concluding remarks in Section V.
@@ II Related Work
Recent research in VANETs and their integration into ITS has significantly focused on improving road safety and efficiency. This section reflects on key advancements in this domain, setting the stage for our novel cyber-twin framework.
@@
Alhaidari et al. introduced a simulation method to create the VANET DDoS dataset, encompassing critical VANET features like traffic density and node mobility. However, it remains inaccessible to the public [6]. Bangui et al. tackled VANETs’ vulnerability to malicious jamming, proposing a data prioritization model enhancing Big Data Analytics’ efficiency in jamming detection, a leap forward in real-time anti-jamming applications [7]. Mokhtar et al. surveyed VANETs’ security, highlighting the challenges due to their mobile, infrastructure-less nature [2]. They emphasized the network’s inherent security vulnerabilities, paving the way for innovative solutions. Another notable study introduced a machine learning-based system for real-time identification of malicious nodes in VANETs, achieving high accuracy with Random Forest (RF) and Gradient-boosted Trees (GBT) models [8].
Figure 2: The state diagram
@@
Haydari et al. developed advanced machine learning methods for intrusion detection in VANETs, targeting sophisticated attacks like DDoS [9]. Their focus on RSUs brought forward a non-parametric IDS, though its effectiveness against sporadic attacks remains a potential area for improvement. Zhou et al. concentrated on enhancing VANET security within ITS for 6G systems, introducing a framework combining identity-based encryption and deep learning, primarily addressing DDoS attack scenarios [10]. Alrehan et al. reviewed VANET architectures, advocating for security solutions uniquely tailored to VANET characteristics, diverging from conventional methods [11]. Anyanwu et al. focused on software-defined network (SDN) integrated VANETs’ vulnerability to DDoS attacks led to an Intrusion Detection Model utilizing a support vector machine classifier, a significant step in vehicle security [12]. A novel self-powered, fog-based VANET infrastructure was also proposed, focusing on sustainable and secure communication, including innovative power management and fog clustering strategies for enhanced network resilience [13].
@@
Within VANET research, the application of digital twin technology remains a relatively underexplored area. Among the few, Arya et al. utilized it for identifying malicious nodes, employing machine learning for practical traffic analysis [14]. Ak et al. developed T6CONF, a new framework that enhances IoT communication in smart cities, thereby improving waste management predictions and achieving net-zero waste objectives [15]. Additionally, Hu et al. explored the Internet of Vehicles (IoV), forming a digital twin connection between physical and virtual vehicles for real-time traffic prediction, highlighting the potential and challenges in handling large, sparse data sets [16].
@@
These studies underscore the untapped potential of VANETs for road safety and network security. However, there is a noticeable gap in effectively combining digital twin technology with AI models for enhanced security and environmental sustainability in VANETs. Our proposed cyber-twin framework aims to bridge this gap, presenting a novel approach to secure vehicle-to-infrastructure communication and advancing green communication in smart city infrastructures.
III Proposed Cyber-twin System Model for Vehicle-to-Infrastructure Security
@@
Our proposed system is designed to foster a resilient and robust communication framework for vehicles in 6G smart cities, leveraging digital twin technology to optimize the computational efficiency of roadside units (RSUs). This enhances the reliability and stability of the RSUs and significantly contributes to green communication initiatives. Some potential benefits of our system can be:
@@
  • Enhanced energy efficiency in RSUs, achieved through reduced computational demands.
@@
Our system adopts a holistic approach with a three-tiered structure comprising physical-to-virtual and virtual-to-virtual communications between layers. This structure forms a comprehensive cyber-twin framework within vehicular ad hoc networks (VANET). Our system is segmented into three layers: data, twinning, and cyber-security, each crucial in ensuring the network’s functionality and resilience, as seen in Fig. 2.
@@ III-B Twinning Layer
This layer is pivotal in creating dynamic digital replicas of VANET’s physical components. It utilizes advanced digital twin modelling and queuing theory to replicate the VANET environment. We have implemented the M/M/m queuing model with a First-In, First-Out strategy for efficient data network simulation. The model evaluates the system’s state by analyzing communication requests against channel availability, providing insights into average waiting times and queue lengths [17].
In this model, channels are likened to servers, and the system’s state is evaluated by comparing the total communication requests against available channels. For the M/M/m queuing model, our key assumptions and parameters are as follows:
@@
  • $\lambda_{R}$ as the arrival rate of communication requests.
  • $\mu$, the channel’s service rate per request, time-dependent.
  • $c_{m}$, the total number of channels.
  • $c_{d}$, the total number of vehicles.
@@
We formulate basic equations using a state diagram in Fig. 2. $c$ represent channel states, $m$ represent the number of channels and $d$ denote the number of vehicles with communication requests. When the number of requests $c_{d}$ is less than or equal to the number of available channels $c_{m}$ (i.e., $c_{d}\leq c_{m}$), all requests are accommodated immediately. However, if the requests exceed the available channels (i.e., $c_{d}>c_{m}$), the system processes requests up to the limit of $c_{m}$, and the remaining requests are queued for subsequent processing. There are $c_{m}$ channels so the probability is increasing ($\mu,2\mu,3\mu\;...\;c_{m}\mu$) until reaching $c_{m}{th}$ channel. After the $c_{m}th$ channel, the leaving of the one request stays the same with probability $c_{m}th$.
@@
Data from the twinning layer is then transferred to the cyber-security layer using a YANG model for robust data transfer.
III-C Cyber-security Layer
The core of our system’s security mechanism is derived from our previous research [18]. It comprises an AutoFS Module and a Multilayer Perceptron, both vital for efficient attack detection.
@@ III-C1 AutoFS Module
III-C2 Labeling Algorithm
Our approach integrates an ensemble learning strategy, merging the Expectation-Maximization (EM) and K-means algorithms to label data effectively. This combination enhances the precision and efficiency of our data categorization process. This method virtually clusters and labels data, considering the infrequency of the attacks for precise prediction.
III-C3 Multilayer Perceptron (MLP)
Our MLP network is optimized for online learning, adapting to continuously changing network data. Its architectural features are a five-layer architecture with relu and softmax activation functions for effective data classification. It employs online learning, continually updating model weights in response to changing network data, ensuring stable performance in a dynamic network environment.
The proposed system model offers a comprehensive and adaptive approach to vehicular communication, which is crucial for the evolving landscape of smart city infrastructure. Its multi-layered architecture bolsters security and optimizes overall performance, positioning it as a pioneering approach to securing 6G smart cities.
@@ IV Performance Evaluation
To test the performance of our proposed system, we used OMNeT++ version 5.1, INET version 3.6, Veins version 4.7, and SUMO version 0.30.0 as the previous work [6]. Our evaluation focused on two main aspects: the efficiency of the digital twin in RSU resource utilization and the effectiveness of the cyber-twin framework in attack detection.
Firstly, we test the performance of the digital twin to the Road Side Unite (RSU). Therefore, we check the average resource usage of the system with the digital twin and without the digital twin. As seen in Fig. 3, when we employed the cyber-twin in the system, the system utilised the resource more efficiently since we reduced the load of the roadside unit thanks to the digital twin. Integrating the cyber-twin system model resulted in more efficient resource utilization, thereby lightening the RSU’s load.
@@
After testing the digital twin effect of the roadside unit resource performance, we experiment with the attack detection performance of the cyber-twin framework. To this end, we used two datasets. The first dataset is the RF Jamming Dataset, which comprises diverse scenarios of RF jamming attacks and interference in VANET [19]. This dataset includes two datasets for two maximum estimated relative speed types. We combined them, adding a new column with the maximum estimated relative speed. The ToN-IoT dataset, specifically designed for assessing the reliability and performance of various AI-driven cybersecurity applications, is utilized in our study [20], since Gad et al. used this dataset for attack detection in VANET. This dataset is particularly tailored for next-generation IoT and industrial IoT scenarios. However, the current version of the ToN-IoT dataset is not VANET-oriented; therefore, we combine the first dataset’s no-attack samples and the ToN-IoT dataset’s attack samples. We selectively used a specified number of samples from these datasets to construct Dataset 2. The details of this selection, including the number of samples from each dataset, are meticulously outlined in Table I.
@@
We compare our solution (PS) with KNN and SVM algorithms since they gave the best result in the following works: [22] and [23], respectively. Fig. 4 and Fig. 5 show the performance comparison of methods for the datasets. As seen in the figures, our solution performs better than the others.
Figure 4: The performance comparison of methods for the dataset 1.
Figure 5: The performance comparison of methods for the dataset 2.
@@
In addition to its robust security and efficiency features, our proposed system significantly contributes to green communication practices in 6G smart city infrastructures. The system’s advanced digital twin technology optimises RSU performance and reduces energy consumption. By efficiently managing data traffic and processing demands, the system minimizes the environmental impact of urban communication networks. This reduction in energy requirements aligns with eco-friendly objectives and diminishes operational costs, making it a sustainable solution for future smart cities. Furthermore, the extended lifespan of hardware components due to reduced strain translates into fewer replacements and maintenance needs, further reinforcing the system’s environmentally conscious design. The cumulative effect of these factors underscores our system’s role in advancing green communication technology, a critical consideration in the development of sustainable urban environments.

### 15 Mar 2024 v4

@@ Abstract
The rapid evolution of Vehicular Ad-hoc NETworks (VANETs) has ushered in a transformative era for intelligent transportation systems (ITS), significantly enhancing road safety and vehicular communication. However, the intricate and dynamic nature of VANETs presents formidable challenges, particularly in vehicle-to-infrastructure (V2I) communications. Roadside Units (RSUs), integral components of VANETs, are increasingly susceptible to cyberattacks, such as jamming and distributed denial of service (DDoS) attacks. These vulnerabilities pose grave risks to road safety, potentially leading to traffic congestion and vehicle malfunctions. Existing methods face difficulties in detecting dynamic attacks and integrating digital twin technology and artificial intelligence (AI) models to enhance VANET cybersecurity. Our study proposes a novel framework that combines digital twin technology with AI to enhance the security of RSUs in VANETs and address this gap. This framework enables real-time monitoring and efficient threat detection while also improving computational efficiency and reducing data transmission delay for increased energy efficiency and hardware durability. Our framework outperforms existing solutions in resource management and attack detection. It reduces RSU load and data transmission delay while achieving an optimal balance between resource consumption and high attack detection effectiveness. This highlights our commitment to secure and sustainable vehicular communication systems for smart cities.
@@ I Introduction
In recent years, the advent of Vehicular Ad-hoc NETworks (VANETs) has marked a significant milestone in the evolution of intelligent transportation systems (ITS). These networks, which seamlessly integrate mobile ad hoc networks (MANET), Internet of Things (IoT), and ITS, are at the forefront of revolutionizing road safety and vehicular communication. The critical role of VANETs in reducing traffic accidents and enhancing road safety cannot be overstated, given the alarming statistics from the World Health Organization highlighting the high incidence of road traffic deaths globally [1].
@@
However, the dynamic and complex nature of VANETs introduces significant challenges, particularly in vehicle-to-infrastructure communications [2]. Road Side Units (RSUs), as critical components of VANETs, are often the targets of cyberattacks, including but not limited to jamming and distributed denial of service (DDoS) attacks, which can severely disrupt vehicular communication and pose significant risks to road safety, leading to potential traffic congestion, vehicle malfunctions, and even accidents [3]. This vulnerability underscores the urgency for innovative solutions to safeguard these systems against such threats. The prevalence of connected and autonomous vehicles in modern urban environments underscores the urgent need to address these vulnerabilities due to the rapid evolution of vehicular technology.
Figure 1: The proposed architecture.
@@
In addressing these challenges, our work focuses primarily on vehicle-to-infrastructure communication, explicitly targeting the security of RSUs. We introduce a novel framework that employs digital twin technology and an advanced artificial intelligence (AI) model. The digital twin concept, a digital replica of a physical asset or system, is central to our approach [4]. By creating a digital twin of RSUs, we achieve a real-time, dynamic representation of the physical units, which allows for the meticulous monitoring and analysis of network interactions. This capability is crucial in detecting and mitigating the aforementioned cyber threats, ensuring the integrity and reliability of vehicle-to-infrastructure communications [5]. We referred to digital twins as ‘cyber-twins’ since we utilized them in the field of cybersecurity. Reducing computational demands and minimizing the average data transmission delay are some of the key areas of research in green VANET technology [6]. Our solution secures communication and contributes to green communications. Our framework focuses on optimizing RSUs’ efficiency by reducing their computational demands and minimizing the average data transmission delay. This aligns with the growing emphasis on sustainable and eco-friendly technologies in smart cities. This focus on both security and sustainability places us at the forefront of ITS advancement. Furthermore, our dedication to creating environmentally friendly technologies for smart cities is highlighted by our efforts to reduce energy consumption and extend the lifespan of hardware. The key contributions of this paper include:
@@
  • We propose a novel framework tailored explicitly for securing vehicle-to-infrastructure communication in VANETs, utilizing cyber twins and an AI model for efficient threat detection.
@@
  • Our approach integrates cyber twins for real-time monitoring, management of vehicular networks, and analysis of RSU performance and security.
@@
  • Proposed framework reduces RSUs’ computational load, improving hardware longevity and energy efficiency.
  • Our solution significantly advances green communications in VANETs by reducing computational demands, minimizing data transmission delays, and contributing to the sustainability of ITS.
@@
Our paper proposes a new approach to enhance vehicle-to-infrastructure communication in VANETs while addressing road safety, sustainability, and security concerns. We use cyber twins and AI models to set a new standard for developing intelligent and secure transportation systems. Our work paves the way for safer and greener smart cities. This work proceeds with an overview of relevant literature in Section II, then delves into the proposed solution in Section III, and evaluates its performance in Section IV. The paper is concluded in Section V.
@@ II Related Work
This section highlights recent advancements in improving road safety and efficiency in VANETs and their integration into ITS, setting the stage for our novel framework.
@@
Alhaidari et al. introduced a simulation method to create the VANET DDoS dataset, encompassing critical VANET features like traffic density and node mobility. However, it remains inaccessible to the public [7]. Bangui et al. tackled VANETs’ vulnerability to malicious jamming, proposing a data prioritization model enhancing Big Data Analytics’ efficiency in jamming detection, a leap forward in real-time anti-jamming applications [8]. Mokhtar et al. surveyed VANETs’ security, highlighting the challenges due to their mobile, infrastructure-less nature [2]. They emphasized the network’s inherent security vulnerabilities, paving the way for innovative solutions. Another notable study introduced a machine learning-based system for real-time identification of malicious nodes in VANETs, achieving high accuracy with random forest and gradient-boosted trees models [9].
Figure 2: The state diagram.
@@
Haydari et al. developed advanced machine learning methods for intrusion detection in VANETs, targeting sophisticated attacks like DDoS [10]. Their RSU-based non-parametric detection system has the potential for improvement against sporadic attacks. Zhou et al. concentrated on enhancing VANET security within ITS for 6G systems, introducing a framework combining identity-based encryption and deep learning, primarily addressing DDoS attack scenarios [11]. Alrehan et al. reviewed VANET architectures, advocating for security solutions uniquely tailored to VANET characteristics, diverging from conventional methods [12]. Anyanwu et al. focused on a software-defined network integrated VANETs’ vulnerability to identify DDoS attacks utilizing a support vector machine (SVM) classifier, a significant step in vehicle security [13]. A novel self-powered, fog-based VANET infrastructure was also proposed, focusing on sustainable and secure communication, including innovative power management and fog clustering strategies for enhanced network resilience [14].
@@
In VANET research, the utilization of digital twin technology has not yet been widely explored. Among the few, Arya et al. utilized it for identifying malicious nodes, employing machine learning for practical traffic analysis [15]. Ak et al. developed T6CONF, a new framework that enhances IoT communication in smart cities, thereby improving waste management predictions and achieving net-zero waste objectives [16]. Additionally, Hu et al. explored the internet of vehicles, forming a digital twin connection between physical and virtual vehicles for real-time traffic prediction, highlighting the potential and challenges in handling extensive, sparse data sets [17].
@@
The studies have highlighted VANETs’ immense potential for improving road safety and network security. However, there is a noticeable gap in effectively combining digital twins with AI models to enhance security and environmental sustainability in VANETs. We propose a new framework to secure vehicle-to-infrastructure communication and advance green communication in smart city infrastructures to address this gap.
III Proposed System Model for Vehicle-to-Infrastructure Security
@@
Our proposed system is designed to foster a resilient and robust communication framework for vehicles in 6G smart cities, leveraging cyber twins to optimize the computational efficiency of RSUs. This enhances the reliability and stability of the RSUs. Some potential benefits of our system can be:
@@
  • Enhanced energy efficiency in RSUs, achieved through reduced computational demands and minimizing the average data transmission delay.
@@
Our system adopts a holistic approach with a three-tiered structure comprising physical-to-virtual and virtual-to-virtual communications between layers. This structure forms a comprehensive cyber-twin framework within VANET. Our system is segmented into three layers: data, twinning, and security, each crucial in ensuring the network’s functionality and resilience, as seen in Fig. 1.
@@ III-B Twinning Layer
This layer is pivotal in creating digital replicas of VANET’s physical components. We used queuing theory to model the VANET environment. We have implemented the M/M/m queuing model with a first-in-first-out strategy for efficient data network simulation. The model evaluates the system’s state by analyzing communication requests against channel availability, providing insights into average waiting times and queue lengths [18].
In this model, channels are linked to servers, and the system’s state is evaluated by comparing the total communication requests against available channels. For the M/M/m queuing model, our key assumptions and parameters are as follows:
@@
  • $\lambda_{R}$ denotes the arrival rate of communication requests.
  • $\mu$ expresses the channel’s service rate per request, which is time-dependent.
  • $c$ represents the channel states.
  • $m$ defines the total number of channels.
  • $d$ denotes the total number of vehicles with communication requests.
@@
We formulate basic equations using a state diagram in Fig. 2. When the number of vehicle requests $c_{d}{{}_{i}}$ is less than or equal to the number of available channels $c_{m}$ (i.e., $c_{d}{{}_{i}}\leq c_{m}$), all requests are accommodated immediately. However, if the requests exceed the available channels (i.e., $c_{d}>c_{m}$), the system processes requests up to the limit of $c_{m}$, and the remaining requests are queued for subsequent processing. There are $c_{m}$ channels so the probability is increasing ($\mu,2\mu,3\mu\;...\;c_{m}\mu$) until reaching $c_{m}$th channel. After the $c_{m}$th channel, the leaving of the one request stays the same with probability $c_{m}$th.
@@
Data from the twinning layer is transferred to the security layer via a YANG model for robust data transfer. We used the YANG model from previous work [19] and adapted it to VANET.
III-C Security Layer
The system’s security mechanism is based on our previous research [20]. It includes an AutoFS Module and a Multilayer Perceptron, both essential for effective attack detection.
@@ III-C1 AutoFS Module
III-C2 Labelling Algorithm
Our labelling method efficiently categorizes data using the expectation-maximization and K-means algorithms, improving precision. The method clusters and labels data to accurately predict attacks, considering their infrequency.
III-C3 Multilayer Perceptron
Our MLP network adapts to continuously changing network data through online learning. It uses a five-layer architecture with relu and softmax activation functions for effective data classification. It constantly updates model weights, ensuring stable performance in a dynamic network environment.
The proposed system model is an adaptive approach to vehicular communication that bolsters security and optimizes performance for smart city infrastructure. It is a pioneering approach to green and secure 6G smart cities.
@@ IV Performance Evaluation
To test the performance of our proposed system, we used OMNeT++ version 5.1, INET version 3.6, Veins version 4.7, and SUMO version 0.30.0 as the previous work [7]. We utilized Eclipse Ditto, which is a scalable and flexible open-source framework, to create cyber twins of physical entities [21]. Our evaluation focused on two main aspects: the efficiency of cyber twins in RSU resource utilization and the effectiveness of our framework in attack detection.
Firstly, we tested the performance of cyber twins in RSU. Therefore, we check the average resource usage of the system with and without the cyber twin. As seen in Fig. 3, when we employed the cyber-twin in the system, the system utilised the resource more efficiently since we reduced the load of the roadside unit. Integrating the cyber-twin system model resulted in more efficient resource utilization, thereby lightening the RSU’s load.
@@
After testing the cyber twin effect of the roadside unit resource performance, we experiment with the attack detection performance of the cyber-twin framework. To this end, we used two datasets. The first dataset is the RF Jamming Dataset [22], which comprises diverse scenarios of RF jamming attacks and interference in VANET. This dataset includes two datasets for two maximum estimated relative speed types. We combined them, adding a new column with the maximum estimated relative speed. The ToN-IoT dataset [23], specifically designed for assessing the reliability and performance of various AI-driven cybersecurity applications, is utilized in our study since Gad et al. used this dataset for attack detection in VANET [24]. This dataset is particularly tailored for next-generation IoT and industrial IoT scenarios. However, the current version of the ToN-IoT dataset is not VANET-oriented; therefore, we combine the first dataset’s no-attack samples and the ToN-IoT dataset’s attack samples. We selectively used a specified number of samples from these datasets to construct Dataset 2. The details of this selection, including the number of samples from each dataset, are meticulously outlined in Table I.
@@
We compare our solution (PS) with KNN and SVM algorithms since they gave the best result in the following works: [25] and [26], respectively. Table II compares the performance of different methods on the datasets. Our solution outperforms the others.
TABLE II: The performance comparison of methods
Precision
F-Measure
Sensitivity
Dataset-1 KNN 95.72 96.28 96.85
SVM 89.27 91.84 94.58
PS 98.96 98.99 99.03
Dataset-2 KNN 88.67 88.90 89.14
SVM 84.67 85.91 87.19
PS 97.53 98.08 98.64
After that, we examined the performance of our solution in terms of average delay and delivery rate under various data message times.
Fig. 4 indicates that PS is performing significantly better in reducing delay than KNN, approximately 1.41 times, and approximately 1.61 times faster than SVM. In terms of average delivery rate, our solution delivers a stable performance. It is approximately 1.11 times more effective than KNN and about 1.21 times more effective than SVM. As can be seen in Fig. 4, PS has significantly lower delays and higher delivery rates compared to the other two methods.
Figure 4: The comparison of methods regarding the average delay and delivery rate under different data message lifetime.
@@
In addition to its robust security and efficiency features, our proposed system significantly contributes to green communication practices in 6G smart city infrastructures. The system’s advanced cyber twins optimise RSU performance and reduce energy consumption. Moreover, it contributes to green communications by reducing their computational demands and minimizing the average data transmission delay. By efficiently managing data traffic and processing demands, the system minimizes the environmental impact of urban communication networks. This reduction in energy requirements aligns with eco-friendly objectives and diminishes operational costs, making it a sustainable solution for future smart cities. Furthermore, the extended lifespan of hardware components due to reduced strain translates into fewer replacements and maintenance needs, further reinforcing the system’s environmentally conscious design. The cumulative effect of these factors underscores our system’s role in advancing green communication technology, a critical consideration in the development of sustainable urban environments.